# The Landscape of College-level Data Visualization Courses, and the Benefits of Incorporating Statistical Thinking

### v1

@@ 3.2 Principle 1: Graphs should be complemented with inference
(a) $n=40$, 95% CIs.
(b) $n=120$, 95% CIs.
(c) $n=200$, 95% CIs.
(d) $n=40$, 98.3% CIs.
(e) $n=120$, 98.3% CIs.
(f) $n=200$, 98.3% CIs.
Figure 5: Bar plots displaying proportions $(0.25,0.35,0.4)$ for a single categorical variable. Left, middle, and right panels correspond to sample sizes $n=40$, 120, and 200, respectively. Plots in the top panels display 95% confidence intervals (CIs), and plots in the bottom panels display Bonferroni-corrected CIs that account for all pairwise comparisons (resulting in 98.3% CIs). Text in upper-left displays $p$-values for a chi-squared test for equal proportions across the three categories.
Indeed, when visually interpreting a graph, it’s natural to make multiple comparisons, such as when we compare confidence intervals. This introduces a multiple testing problem, a complication with exploratory data visualization that has been acknowledged in various literatures (e.g., Buja et al., (2009); Zhao et al., (2017); Zgraggen et al., (2018); Savvides et al., (2022)). The bottom row of Figure 5 displays confidence intervals using a Bonferonni correction for all three paired comparisons; this is equivalent to displaying $100(1-0.05/3)\%=98.3\%$ confidence intervals. This causes more intervals to overlap and changes some testing conclusions for the $n=120$ plot, as shown in Table 2, which summarizes inferential conclusions that arise from the six bar plots in Figure 5. Walking through this example with students demonstrates how they can correct for multiple testing not just in statistical analyses, but also in visualizations.
@@
At first, many students think bar plots are too basic for a college-level data visualization course. Indeed, a bar plot is perhaps the simplest graph one could imagine; it’s nonetheless a rich source for inferential nuances. This sets the stage for the rest of the semester, where students have to pinpoint which inferential tools are most appropriate for a particular graph. For example, when we teach visualizing quantitative distributions with overlayed densities, we acknowledge that there are many potential ways to compare distributions (e.g., inspecting means, variances, or distributional shapes), which correspond to many potential analyses (e.g., ANOVA, Bartlett’s test, or the KS test). Once again, this introduces a multiple testing problem and issues of statistical power, because comparing shapes will have less power than comparing means. Students can more readily grasp these complications in more nuanced graphs like overlayed densities once they’ve seen a simpler example like Figure 5.
The primary takeaway is that, for almost any graph (simple or complex), students should consider how to determine if a graphical phenomenon is “statistically significant,” and what they can conclude from their graphs and analyses together.
@@ 3.3 Principle 2: Graphs are estimates and statistics
First we discuss the pedagogical value of viewing graphs as estimates. For example, when we teach visualizing and analyzing two quantitative variables with scatterplots and linear regression, we emphasize that the regression line drawn on the scatterplot is indeed an estimator that has a distribution, which motivates displaying confidence bands for that line. One way to interpret 95% confidence bands for a regression line is: If we obtained many random samples from the same population, and visualized each sample with a scatterplot and confidence bands, then approximately 95% of those visual bands will contain the true line. [^6] Technically, this interpretation is only true when one displays uniform confidence bands, rather than pointwise confidence bands. This inferential nuance is also a useful discussion point for students. Although our undergraduate students are familiar with the definition of confidence intervals, this interpretation is often novel for them, because it suggests a population-level graph and a distribution for that graph among samples.
@@
To illustrate a “distribution of graphs,” we’ll consider visualizing a quantitative measure of loudness (in decibels) among songs on Spotify, averaged within years from 1922-2021.777The original dataset is available on Kaggle at https://www.kaggle.com/datasets/yamaerenay/spotify-dataset-19212020-600k-tracks. Thus, the dataset consists of 100 observations, corresponding to the “average” song for a given year. Let’s consider visualizing the distribution of loudness with a smoothed density, which involves estimating a bandwidth to quantify the distribution’s smoothness. For example, the density() function in R uses Silverman’s rule-of-thumb (Silverman,, 2018) by default for bandwidth estimation. This would lead to one smoothed density plot, but because the bandwidth is estimated, there is uncertainty in the distribution’s smoothness and the plot itself. To understand this uncertainty, we created 1000 bootstrapped datasets via sampling with replacement, and computed the estimated bandwidth via Silverman’s rule-of-thumb for each bootstrapped dataset. Then, we computed the average bandwidth, $\bar{h}$, across these 1000 bootstrapped datasets, as well as 2.5% and 97.5% quantiles, denoted $h_{(0.025)}$ and $h_{(0.975)}$. Figure 6 displays three smoothed densities, corresponding to using $\bar{h}$, $h_{(0.025)}$, or $h_{(0.975)}$ as the bandwidth. The density using the average bandwidth $\bar{h}$ (in black) clearly has two modes on the right and middle of the graph, but it’s less clear if there is a smaller third mode on the left. This smaller mode is more distinct when we use the smaller bandwidth $h_{(0.025)}$, but it’s virtually nonexistent when we use the larger bandwidth $h_{(0.975)}$; meanwhile, the other two modes are clearly visible in both plots. This suggests that we have some uncertainty about whether this is a bimodal or trimodal distribution. More generally, Figure 6 visualizes a “confidence set” of graphs we may reasonably observe across random samples.
@@
Figure 6: Distribution of loudness of songs, averaged within years from 1922-2021. Using this dataset of 100 observations, we generated 1000 bootstrapped datasets, and computed the estimated bandwidth for each bootstrapped dataset. Displayed is the density according to the average bandwidth (solid black), and the densities according to the lower 2.5% quantile of bandwidths (dashed red) and upper 97.5% quantile bandwidth (dashed blue).
@@
From a theoretical standpoint, this bootstrapping approach is not necessarily the optimal way to conduct inference for graphical properties like a distribution’s smoothness and modes. For example, there is a rich literature on how to construct confidence sets for smoothed densities (e.g., Wasserman, 2006, Chapter 6) and conduct inference for modes in a distribution (e.g., Genovese et al., 2016). Nonetheless, a bootstrap procedure is a useful way to demonstrate to students that graphs indeed have a distribution, because they are estimators. For example, Hesterberg, (2015) discusses the pedagogical benefits of using a “graphical bootstrap” to illustrate the uncertainty of linear regression lines. Furthermore, bootstrapping procedures are relatively easy to implement and can flexibly incorporate many graphs, such that they have been used in recent statistical software to quantify uncertainty in data visualizations (Waskom,, 2021; Sun et al.,, 2023). For students, the important takeaway is that a graph is a statistic, because it is a function of observed data, just like a sample mean or median. Graphs can thus act as estimates for population-level quantities.
@@
Because graphs are statistics, they can also be used as test statistics within a hypothesis testing procedure. In short, we can consider the “null distribution” of a graph under a given hypothesis, and view the graph we create for a given dataset as the observed test statistic. Using this insight, many works have proposed a “lineup protocol,” where the observed graph is plotted alongside graphs under a null distribution, and a test is conducted by determining if a user can differentiate the observed graph from graphs under the null (Buja et al.,, 2009; Wickham et al.,, 2010; Majumder et al.,, 2013; Vanderplas et al.,, 2020). Buja et al., (2009) discuss many examples of this procedure, such as testing temporal dependence by permuting and plotting many time series graphs. Meanwhile, Majumder et al., (2013) studied this procedure within the context of linear regression hypotheses (e.g., that a coefficient is equal to zero), and conducted several real experiments with human subjects to validate that this procedure has high statistical power. As an illustrative example, we consider the “lineup” procedure for maps (Walter,, 1993; Wickham et al.,, 2010), to demonstrate how to apply this procedure to complex graphics, as we teach in the second half of our course.
@@
Figure 7 illustrates how the lineup procedure operates for areal spatial data using choropleths, where geographic regions are colored to denote variable values. Using data from the 2015 American Community Survey, we plotted the average unemployment rate in each of the 48 contiguous United States. [^8] The original dataset is available on Kaggle at https://www.kaggle.com/datasets/census/2015-american-community-survey. Variables are measured at the county level. To compute averages at the state level, we computed a weighted average, where weights corresponded to county population size. Using the resulting map, we may ask: Is there spatial dependence in unemployment rates? We can answer this question from a testing perspective, where we consider what the distribution of this map is under the null hypothesis of spatial independence. To generate maps under this null hypothesis, we permuted the data across states and created a map with the resulting permuted data. Figure 7 displays 24 of these maps alongside the real map; one conducts the hypothesis test by selecting the map that appears the most distinct among the 25 maps. Thus, if one selects the real map, the $p$-value is $1/25=0.04$, such that one would reject the null hypothesis of spatial independence at the $\alpha=0.05$ level. The identity of the real map is revealed in the replication materials. Students are often surprised by the amount of apparent spatial clustering there can be in random maps; this gives them appreciation for the phenomenon that humans often find patterns in seemingly random visual stimuli, known as “pareidolia.”
@@
Figure 7: Average unemployment rate in each of the 48 contiguous United States, according to the 2015 American Community Survey. Displayed are 24 “random maps” where the data is randomly permuted across states, alongside the real map. The identity of the real map is revealed in the replication materials. The color coding is set such that white denotes the median unemployment rate across states.
@@ 3.4 Principle 3: Graphs can be used to motivate, teach, and interpret statistical analyses
This study thus has two categorical variables and two quantitative variables. As is common in data visualization courses, we teach students how to display all four variables in a single scatterplot, where quantitative variables are displayed on the x- and y-axes and categorical variables are denoted by color and shape. The color and shape are known as “aesthetics” in the grammar of graphics framework, in the sense that data can be mapped to visual aesthetics (Wilkinson,, 2012). Regression lines are also an aesthetic; thus, scatterplots can implicitly display different regression models, depending on how variables are aesthetically mapped to regression lines.
@@
Figure 8 shows four scatterplots that display unemployment (x-axis) and voter turnout (y-axis), and use color and shape to denote the female and African American variables, respectively. The only difference among the plots is how color and shape are mapped to regression lines: No aesthetics are mapped to the regression in Plot 1 (resulting in one regression line), only color or shape are mapped in Plots 2 and 3 (resulting in two lines), and both color and shape are mapped in Plot 4 (resulting in four lines). Students often find these visual differences interesting because they correspond to subtle variations in code that many students tend to write. Students can readily see that the plots in Figure 8 display regression models for different subsets of the data, based on the female and African American variables. We then have students implement the regression models implicitly displayed by each plot. For example, Table 3 shows the estimated coefficients for a model that regresses voter turnout on unemployment, the female and African American variables, and their interactions, which corresponds to the regression displayed in Plot 4. Given this plot, students can readily interpret the estimated coefficients in terms of changes in intercepts and slopes—an insight that’s often difficult for students to grasp if they only look at results like those in Table 3.
@@
There are two intended lessons from this example. First, there is often an implicit statistical model suggested by a graph, depending on how variables are mapped to the graph’s aesthetics. It can be useful to have students consider which model is suggested by a particular graph, and then implement that model; this lets the graph motivate the analysis, which is often intuitive for students. Second, by placing graphs and their corresponding analyses side-by-side, students can more quickly understand and interpret statistical analyses by seeing how they operate graphically. This also brings students closer to being effective data scientists, who use graphs and analyses synergistically to arrive at comprehensive conclusions.
@@
Figure 8: Scatterplots of municipal-level unemployment rates and voter turnout in Louisiana mayoral elections between 1988-2011. Points are colored and shaped by whether there was at least one female candidate and at least one African American candidate in a given election, respectively. The only difference among the plots is how points’ color and shape are mapped to regression lines, which in turn corresponds to different statistical models. In Plot 1, no aesthetics are mapped to the regression line. In Plots 2 and 3, only the shape or color, respectively, are mapped to the regression lines. In Plot 4, both shape and color are mapped to the regression lines.
@@
A biplot (Gabriel,, 1971) is particularly useful for this task, which in its simplest form overlays the scatterplot with arrows visualizing these loadings, as displayed in Figure 9(a). There is an arrow for each variable; the length and direction of the arrows denote the magnitude and sign, respectively, of the loadings, thereby visualizing the variables’ relationship with these two principal components. For example, we can see that more recent songs have a higher tempo, energy, loudness, and popularity, whereas older songs have higher levels of “instrumentalness” and “acousticness.” This helps students gain deeper insights about visualizations (e.g., how to interpret scatterplots of principal components) as well as analyses (e.g., understanding what the loadings communicate substantively).
Meanwhile, Figure 9(b) shows a scree plot for the Spotify data. It’s somewhat subjective where the “elbow” in the plot is; one can argue that it is anywhere between $k=4$ and $k=7$ components. Instead of only focusing on the “elbow,” we ask students to consider what the “variance explained” (y-axis) means in terms of scatterplots of components. Within a scatterplot, a component’s variance corresponds to horizontal or vertical spread. Thus, plotting the first two components corresponds to the scatterplot with the most horizontal and vertical spread—in this case, $57.5\%+18.8\%=76.3\%$ of the total spread. If we decided to use additional components (e.g. based on the scree plot figure), we would make additional scatterplots, but those scatterplots would have much less visual spread than the one in Figure 9(a). This helps students understand the concept of “variance explained” within PCA in terms of visual spread, and why organizing components by variance is useful for visualization.
@@
(a) Biplot of the top two principal components and 11 variables. Color denotes decades that songs were released. Length and direction of arrows reflect variables’ relationship with the principal components. Note that the variable labels “popularity” and “loudness” overlap in the upper-right.
(b) Scree plot. Variance explained (y-axis) is defined as each principal component’s variance divided by the sum of variances (or equivalently, the number of principal components). One can argue that the “elbow” is anywhere between four and seven principal components.

### v2

@@ 3.2 Principle 1: Graphs should be complemented with inference
Figure 5: Bar plots displaying proportions $(0.25,0.35,0.4)$ for a single categorical variable; each y-axis is on the proportion scale. Plots in the left, middle, and right columns correspond to sample sizes $n=40$, 120, and 200. Plots in the top row display 95% confidence intervals (CIs, in black), and plots in the bottom row display Bonferroni-corrected CIs that account for all pairwise comparisons (in orange). The title of each plot also lists the $p$-value for a chi-squared test for equal proportions across the three categories.
Indeed, when visually interpreting a graph, it’s natural to make multiple comparisons, such as when we compare confidence intervals. This introduces a multiple testing problem, because we are visually making many comparisons within a single graph (Wright et al., , 2019). The bottom row of Figure 5 displays confidence intervals using a Bonferonni correction for all three paired comparisons; this is equivalent to displaying $100(1-0.05/3)\%=98.3\%$ confidence intervals. This causes more intervals to overlap and changes some testing conclusions for the $n=120$ plot, as shown in Table 2, which summarizes inferential conclusions that arise from the six bar plots in Figure 5. Walking through this example with students demonstrates how they can correct for multiple testing not just in statistical analyses but also in visualizations.
@@
It’s worth noting that it’s often critiqued to use bar plots with confidence intervals to display uncertainty (Newman and Scholl, , 2012; Weissgerber et al., , 2015); such visuals are sometimes referred to as “dynamite plots” or “plunger plots.” Dynamite plots use bars to display sample means and confidence intervals of a quantitative variable across different categories. Such plots are critiqued because, instead of displaying only sample means of a quantitative variable, it is more informative to visualize its distribution (e.g., via histograms, boxplots, or smoothed densities, as we note in the “1D Quantitative & 1D Categorical” row in Table 1). Meanwhile, Figure 5 displays just a single categorical variable: The height of the bars denotes the proportion in each category, rather than the sample mean of a quantitative variable, and thus would not be considered a dynamite plot.
Indeed, because bar plots visualize just a single categorical variable, many students may think that they are too basic for a college-level data visualization course. Figure 5 demonstrates that many inferential nuances nonetheless arise even in this simple setting. This sets the stage for the rest of the semester, where students have to pinpoint which inferential tools are most appropriate for a particular graph. For example, when we teach visualizing quantitative distributions with overlayed densities, we acknowledge that there are many potential ways to compare distributions (e.g., inspecting means, variances, or distributional shapes), which correspond to many potential analyses (e.g., ANOVA, Bartlett’s test, or the KS test). Once again, this introduces a multiple testing problem and issues of statistical power, because comparing shapes will have less power than comparing means. Students can more readily grasp these complications in more nuanced graphs like overlayed densities once they’ve seen a simpler example like Figure 5. The primary takeaway is that, for almost any graph (simple or complex), students should consider how to determine if a graphical phenomenon is “statistically significant,” and what they can conclude from their graphs and analyses together.
Our discussion above focuses on inferential complications that can arise when examining a single graph. However, even more complications arise when we consider graphs and analyses made in sequence. For example, perhaps we conclude that population means of a quantitative variable differ across categories, based on a side-by-side boxplot and one-way ANOVA test; then, we may decide to create another graph and analysis to conduct another comparison. As a result, inference based on the second graph should be conditional on inference from the first graph. This is related to the literature on post-selection inference (e.g., Taylor and Tibshirani, 2015; Tibshirani et al., 2016), where one tries to conduct inference for a regression model conditional on certain variables being selected for that model (e.g., via stepwise selection)—but in this case, we must condition on results from previous graphs, instead of from selection procedures. Although this complication in exploratory data analyses has been acknowledged in various literatures (Buja et al., , 2009; Zhao et al., , 2017; Zgraggen et al., , 2018; Savvides et al., , 2022), we do not discuss this in-depth in our class, in part because it is still an open problem how to validly conduct inference conditional on a sequence of visualizations. Nonetheless, grappling with this complication requires conceptualizing graphs as estimates and statistics that are used for inference—an idea we discuss next.
@@ 3.3 Principle 2: Graphs are estimates and statistics
First we discuss the pedagogical value of viewing graphs as estimates. For example, when we teach visualizing and analyzing two quantitative variables with scatterplots and linear regression, we emphasize that the regression line drawn on the scatterplot is indeed an estimator that has a distribution, which motivates displaying confidence bands for that line. One way to interpret 95% confidence bands for a regression line is: If we obtained many random samples from the same population, and visualized each sample with a scatterplot and confidence bands, then approximately 95% of those visual bands will contain the true line. [^6] Technically, this interpretation is only true when one displays uniform confidence bands, rather than pointwise confidence bands. This inferential nuance is also a useful discussion point for students. Although our undergraduate students are familiar with the definition of confidence intervals, this interpretation is often novel for them, because it suggests a population-level graph and variation of a graph from sample to sample.
@@
To illustrate how graphs can vary from sample to sample, we’ll consider visualizing a quantitative measure of loudness (in decibels) among songs on Spotify, averaged within years from 1922-2021.777The original dataset is available on Kaggle at https://www.kaggle.com/datasets/yamaerenay/spotify-dataset-19212020-600k-tracks; the dataset averaged across years is available in our replication materials. Thus, the dataset consists of 100 observations, corresponding to the “average” song for a given year. Let’s consider visualizing the distribution of loudness with a smoothed density, which involves estimating a bandwidth to quantify the distribution’s smoothness. For example, the density() function in R uses Silverman’s rule-of-thumb (Silverman, , 2018) by default for bandwidth estimation. This would lead to one smoothed density plot, but because the bandwidth is estimated, there is uncertainty in the distribution’s smoothness and thus the plot itself. To understand this uncertainty, we created 1000 bootstrapped datasets via sampling with replacement, and computed the estimated bandwidth via Silverman’s rule-of-thumb for each bootstrapped dataset. Then, we computed the average bandwidth, $\bar{h}$, across these 1000 bootstrapped datasets, as well as 2.5% and 97.5% quantiles, denoted $h_{(0.025)}$ and $h_{(0.975)}$. Figure 6 displays three smoothed densities, corresponding to using $\bar{h}$, $h_{(0.025)}$, or $h_{(0.975)}$ as the bandwidth. The density using the average bandwidth $\bar{h}$ (in black) clearly has two modes on the right and middle of the graph, but it’s less clear if there is a smaller third mode on the left. This smaller mode is more distinct when we use the smaller bandwidth $h_{(0.025)}$, but it’s virtually nonexistent when we use the larger bandwidth $h_{(0.975)}$; meanwhile, the other two modes are clearly visible in both plots. This suggests that we have some uncertainty about whether this is a bimodal or trimodal distribution.
@@
Figure 6: Distribution of loudness of songs, averaged within years from 1922-2021. Using this dataset of 100 observations, we generated 1000 bootstrapped datasets, and computed the estimated bandwidth for each bootstrapped dataset. Displayed is the density according to the average bandwidth (solid black), and the densities according to the lower 2.5% quantile of bandwidths (dashed orange) and upper 97.5% quantile bandwidth (dashed blue).
@@
From a theoretical standpoint, this bootstrapping approach is not necessarily the optimal way to conduct inference for graphical properties like a distribution’s smoothness and modes. For example, there is a rich literature on how to construct confidence sets for smoothed densities (e.g., Wasserman, 2006, Chapter 6) and conduct inference for modes in a distribution (e.g., Genovese et al., 2016). Furthermore, it’s important to emphasize that this bootstrapping approach only illustrates variation in graphical properties, like the bandwidth for smoothed densities, rather than the full variation in the graph itself. The reason is that graphical properties are often numeric summaries where averages and quantiles are well-understood, whereas it’s more difficult to consider these operations on visuals. For example, Figure 6 only visualizes three smoothed densities, resulting from the average bandwidth and its quantiles, and does not necessarily visualize the “average” graph or the “quantiles” of such graphs. Nonetheless, a bootstrap procedure is a useful way to demonstrate to students that graphs indeed have sample-to-sample variation. For example, Hesterberg, (2015) discusses the pedagogical benefits of using a “graphical bootstrap” to illustrate the uncertainty of linear regression lines. Furthermore, bootstrapping procedures are relatively easy to implement and can flexibly incorporate many graphs, such that they have been used in recent statistical software to quantify uncertainty in data visualizations (Waskom, , 2021; Sun et al., , 2023). For students, the important takeaway is that graphs are statistics, because they are functions of observed data, just like sample means or medians. Graphs thus act as estimates for population-level quantities.
@@
Because graphs are statistics, they can also be used as test statistics within a hypothesis testing procedure. In short, we can consider the “null distribution” of a graph under a given hypothesis, and view the graph we create for a given dataset as the observed test statistic. Using this insight, many works have proposed a “lineup protocol,” where the observed graph is plotted alongside graphs under a null distribution, and a test is conducted by determining if a user can differentiate the observed graph from graphs under the null (Buja et al., , 2009; Wickham et al., , 2010; Majumder et al., , 2013; Loy et al., , 2016; Vanderplas et al., , 2020; Loy, , 2021; Li et al., , 2024). Buja et al., (2009) and Vanderplas et al., (2020) discuss many examples of this procedure, such as testing temporal dependence by permuting and plotting many time series graphs. Majumder et al., (2013) studied this procedure within the context of linear regression hypotheses (e.g., that a coefficient is equal to zero), and conducted several real experiments with human subjects to validate that this procedure has high statistical power. Loy et al., (2016) and Li et al., (2024) found similar empirical evidence when using the lineup protocol for q-q plots and residual diagnostics, respectively. Meanwhile, Loy, (2021) discuss how lineup protocols are a useful pedagogical tool for teaching students the concepts of signal-versus-noise and variability. As an illustrative example, we consider the lineup protocol for maps (Walter, , 1993; Wickham et al., , 2010), to demonstrate how to apply this procedure to complex graphics, as we teach in the second half of our course.
@@
Figure 7 illustrates how the lineup protocol operates for areal spatial data using choropleths, where geographic regions are colored to denote variable values. Using data from the 2015 American Community Survey, we plotted the average unemployment rate in each of the 48 contiguous United States. [^8] The original dataset is available in our replication materials, as well as on Kaggle at https://www.kaggle.com/datasets/census/2015-american-community-survey. Variables are measured at the county level. To compute averages at the state level, we computed a weighted average, where weights corresponded to county population size. Using the resulting map, we may ask: Is there spatial dependence in unemployment rates? We can answer this question from a testing perspective, where we consider what the distribution of this map is under the null hypothesis of spatial independence. To generate maps under this null hypothesis, we permuted the data across states and created a map with the resulting permuted data. Figure 7 displays 24 of these maps alongside the real map; one conducts the hypothesis test by selecting the map that appears the most distinct among the 25 maps. Thus, if one selects the real map, the $p$-value is $1/25=0.04$, such that one would reject the null hypothesis of spatial independence at the $\alpha=0.05$ level. Students are often surprised by the amount of apparent spatial clustering there can be in random maps; this gives them appreciation for the phenomenon that humans often find patterns in seemingly random visual stimuli, known as pareidolia.
@@
Figure 7: Average unemployment rate in each of the 48 contiguous United States, according to the 2015 American Community Survey. Displayed are 24 “random maps” where the data is randomly permuted across states, alongside the real map (which is in the second row, fifth column). The color coding is set such that white denotes the median unemployment rate across states.
@@ 3.4 Principle 3: Graphs can be used to motivate, teach, and interpret statistical analyses
This study thus has two categorical variables and two quantitative variables. As is common in data visualization courses, we teach students how to display all four variables in a single scatterplot, where quantitative variables are displayed on the x- and y-axes and categorical variables are denoted by color and shape. Color and shape are known as “aesthetics” in the grammar of graphics framework, in the sense that data can be mapped to visual aesthetics (Wickham, , 2010; Wilkinson, , 2012). The following example demonstrates that scatterplots can implicitly display different regression models, depending on how variables are aesthetically mapped to regression lines.
@@
Figure 8 shows four scatterplots that display unemployment (x-axis) and voter turnout (y-axis), and use color and shape to denote the female and African American variables, respectively. The only difference among the plots is how color and shape are mapped to regression lines: No aesthetics are mapped to the regression in Plot 1 (resulting in one regression line), only color or shape are mapped in Plots 2 and 3 (resulting in two lines), and both color and shape are mapped in Plot 4 (resulting in four lines). Students often find these visual differences interesting because they correspond to subtle variations in code that many students tend to write. However, this example demonstrates that once we map aesthetics to statistical models, inferential ambiguities can arise. For example, students can readily see that the plots in Figure 8 display regression models for different subsets of the data: Plot 1 shows a regression line for the entire dataset, and Plot 3 shows an orange regression line for elections with female candidates and a black regression line for elections without female candidates. Meanwhile, Plot 2 shows two regression lines—one for elections with African American candidates and one for those without—but it’s ambiguous which line corresponds to which subset of data, because shapes cannot readily be mapped to lines. A similar issue arises with Plot 4. Given these plots, we ask students to resolve these ambiguities by having them implement the regression models implicitly displayed by each plot. For example, Table 3 shows the estimated coefficients for a model that regresses voter turnout on unemployment, the female and African American variables, and their interactions, which corresponds to the regression displayed in Plot 4. Students can then use this output to determine which regression line corresponds to which subset of data by interpreting the estimated coefficients in terms of changes in intercepts and slopes. For example, because the estimated coefficient for the three-way interaction is negative (and is greater in magnitude than the estimated coefficients for the two-way interactions involving unemployment), we can infer that the downward-sloping orange line corresponds to the regression for elections that have African American and female candidates. We’ve found that students can more readily realize that interactions in linear regression models correspond to changes in intercepts and slopes when they interpret output like Table 3 alongside its corresponding graph.
Grappling with the inferential ambiguities in Figure 8 can also help students appreciate the implications—and potential downsides—of certain visual design choices. Humans tend to focus on differences in color over differences in shape (Wagemans et al., , 2012; Garner, , 2014), such that Figure 8 implicitly emphasizes the orange-black differences over the circle-triangle differences. By pinpointing which statistical model in Figure 8 we’d like to emphasize, we can determine how to improve a plot’s graphic design. For example, if we wanted to utilize a model that only considers differences by the African American variable (as in Plot 2) or by the female variable (as in Plot 3), we may only use color for the variable of interest and keep all observations the same shape. On the other hand, if we wanted to utilize a model that considers the interaction between these variables (as in Plot 4), we may use four colors to denote these four groups, or we may facet by one categorical variable to create two scatterplots that each use color to differentiate observations.
@@
There are three intended lessons from this example. First, there is often an implicit statistical model suggested by a graph, depending on how variables are mapped to the graph’s aesthetics. It can be useful to have students consider which model is suggested by a particular graph, and then implement that model; this lets the graph motivate the analysis, which is often intuitive for students. Second, by placing graphs and their corresponding analyses side-by-side, students can more quickly understand and interpret statistical analyses by seeing how they operate graphically. Third, by assessing whether a graphic aligns with an intended statistical model, students can pinpoint how to improve their graphs’ visual design. This brings students closer to being effective data scientists, who use graphs and analyses synergistically to arrive at comprehensive conclusions.
@@
Figure 8: Scatterplots of municipal-level unemployment rates and voter turnout in Louisiana mayoral elections between 1988-2011. Points are colored and shaped by whether there was at least one female candidate and at least one African American candidate in a given election, respectively. The only difference among the plots is how points’ color and shape are mapped to regression lines, which in turn corresponds to different statistical models. In Plot 1, no aesthetics are mapped to the regression line. In Plots 2 and 3, only the shape or color, respectively, are mapped to regression lines. In Plot 4, both shape and color are mapped to regression lines.
@@
A biplot (Gabriel, , 1971) is particularly useful for this task, which in its simplest form overlays the scatterplot with arrows visualizing these loadings, as displayed in Figure 9(a). There is an arrow for each variable; the length and direction of the arrows denote the magnitude and sign, respectively, of the loadings, thereby visualizing the variables’ relationship with these two principal components. This allows us to interpret observations in the principal component space in terms of the original variables in the data. For example, in Figure 9(a), observations on the left tend to have higher levels of “instrumentalness” and “acousticness,” whereas observations on the right tend to have higher levels of popularity, loudness, energy, and tempo. Because songs become more recent as we go from left to right, this suggests that more recent songs tend to have a higher tempo, energy, loudness, and popularity, while older songs tend to have higher levels of “instrumentalness” and “acousticness.” That said, there may be other unmeasured variabes (besides decade of release) that differentiate observations from left-to-right in the plot. For example, we could have taken an unsupervised approach by applying $k$-means clustering to the observations in this two-dimensional space, and then used the loadings to interpret the resulting clusters, without reference to decade of release. In any case, the biplot can help students gain deeper insights about visualizations (e.g., how to interpret scatterplots of principal components) as well as analyses (e.g., understanding what the loadings communicate substantively).
The biplot in Figure 9(a) only partially captures potential clustering of observations, because it’s projecting a 10-dimensional space down to two dimensions. This is where a scree plot can be helpful for understanding how much variation in the data we’re ignoring when interpreting the biplot. Figure 9(b) shows a scree plot for the Spotify data. It’s somewhat subjective where the “elbow” in the plot is; one can argue that it is anywhere between $k=4$ and $k=7$ components. Instead of only focusing on the “elbow,” we ask students to consider what the “variance explained” (y-axis) means in terms of scatterplots of components. Within a scatterplot, a component’s variance corresponds to horizontal or vertical spread. Thus, plotting the first two components corresponds to the scatterplot with the most horizontal and vertical spread—in this case, $57.5\%+18.8\%=76.3\%$ of the total spread. If we decided to use additional components, we would make additional scatterplots, but those scatterplots would have much less visual spread than the one in Figure 9(a). That said, Figure 9(a) ignores 23.7% of the variation in the data, which could reveal other potential clusters if we examined it with further visualizations. This helps students understand the concept of “variance explained” within PCA in terms of visual spread, and why organizing components by variance is useful for visualization.
@@
(a) Biplot of the top two principal components and 11 variables. Color denotes decades that songs were released. Length and direction of arrows reflect variables’ relationship with the principal components.
(b) Scree plot. Variance explained (y-axis) is defined as each principal component’s variance divided by the sum of variances. One can argue that the “elbow” is anywhere between four and seven principal components.